Find the general solution of the differential equations using method of unndetermined coefficients (MUC) and method of variation of parameters (MVP) .

y" +y = 2x sin x
1
Expert's answer
2021-09-29T07:22:30-0400

1. Homogeneous equation


"y''+y=0"

Corresponding (auxiliary) equation


"r^2+r=0"

"r=\pm i"

The general solution of the homogeneous differential equation is


"y_h=c_1\cos x+c_2\sin x"

Find the particular solution of the non homogeneous differential equation


"y_p=(Ax^2+Bx+C)\cos x+(Dx^2+Ex+F)\sin x"

"y_p'=-(Ax^2+Bx+C)\sin x+(2Ax+B)\cos x"

"+(Dx^2+Ex+F)\cos x+(2Dx+E)\sin x"

"y_p''=-(Ax^2+Bx+C)\cos x-2(2Ax+B)\sin x"

"+2A\cos x-(Dx^2+Ex+F)\sin x"

"+2(2Dx+E)\cos x+2D\sin x"

Substitute


"-(Ax^2+Bx+C)\cos x-2(2Ax+B)\sin x""+2A\cos x-(Dx^2+Ex+F)\sin x""+2(2Dx+E)\cos x+2D\sin x""+(Ax^2+Bx+C)\cos x+(Dx^2+Ex+F)\sin x""=2x\sin x"



"\sin x:"


"-4Ax-2B-Dx^2-Ex-F+2D"

"+Dx^2+Ex+F=2x"

"\cos x:"


"-Ax^2-Bx-C+2A+4Dx+2E"

"+Ax^2+Bx+C=0"

"-4A=2""-2B+2D=0""D=0""2A+2E=0"

"y_p=-\dfrac{1}{2}x^2\cos x+\dfrac{1}{2}x\sin x"

The general solution of the non homogeneous differential equation is


"y=y_h+y_p"

"y=c_1\cos x+c_2\sin x-\dfrac{1}{2}x^2\cos x+\dfrac{1}{2}x\sin x"

2. Homogeneous equation


"y''+y=0"

Corresponding (auxiliary) equation


"r^2+r=0"

"r=\pm i"

Find the general solution of the non homogeneous differential equation in form


"y=C_1\cos x+C_2\sin x"

"y'=-C_1\sin x+C_2\cos x+C_1'\cos x+C_2'\sin x"

Let


"C_1'\cos x+C_2'\sin x=0"

Then


"y'=-C_1\sin x+C_2\cos x"

"y''=-C_1\cos x-C_2\sin x-C_1' \sin x+C_2'\cos x"

Substitute


"-C_1\cos x-C_2\sin x-C_1' \sin x+C_2'\cos x"

"+C_1\cos x+C_2\sin x=2x\sin x"

"C_1'=-C_2'\tan x"

"C_2' \tan x\sin x+C_2'\cos x=2x\sin x"

"C_2' \sin^2 x+C_2'\cos^2 x=2x\sin x\cos x"

"C_2'=2x\sin x\cos x"

"C_2=\int2x\sin x\cos xdx"

"\int2x\sin x\cos xdx"




"\int udv=uv-\int vdu"

"u=x, du=dx"

"dv=2\sin x\cos xdx, v=\int2\sin x\cos xdx"

"=\int\sin (2x)dx=-\dfrac{1}{2}\cos(2x)"

"\int2x\sin x\cos xdx=-\dfrac{1}{2}x\cos(2x)+\int\dfrac{1}{2}\cos(2x)dx"

"=-\dfrac{1}{2}x\cos(2x)+\dfrac{1}{4}\sin(2x)+C_3"

"C_2=-\dfrac{1}{2}x\cos(2x)+\dfrac{1}{4}\sin(2x)+C_3"


"C_1'=-2x\sin x\cos x\tan x=-2x\sin^2x"

"=-x(1-\cos (2x)=x\cos (2x)-x"

"C_1=\int(x\cos (2x)-x)dx"

"\int x\cos (2x)dx"


"\int udv=uv-\int vdu"

"u=x, du=dx"

"dv=\cos (2x)dx, v=\int\cos(2x)dx=\dfrac{1}{2}\sin(2x)"

"\int x\cos (2x)dx=\dfrac{1}{2}x\sin(2x)-\int\dfrac{1}{2}\sin(2x)dx"

"=\dfrac{1}{2}x\sin(2x)+\dfrac{1}{4}\cos(2x)+C_4"

"C_1=\dfrac{1}{2}x\sin(2x)+\dfrac{1}{4}\cos(2x)-\dfrac{1}{2}x^2+C_4"

"y=(\dfrac{1}{2}x\sin(2x)+\dfrac{1}{4}\cos(2x)-\dfrac{1}{2}x^2+C_4)\cos x"

"+(-\dfrac{1}{2}x\cos(2x)+\dfrac{1}{4}\sin(2x)+C_3)\sin x"

"y=C_4\cos x+C_3\sin x-\dfrac{1}{2}x^2\cos x"

"+\dfrac{1}{2}x(\sin(2x)\cos x-\cos(2x)\sin x)"

"+\dfrac{1}{4}(\cos(2x)\cos x+\sin(2x)\sin x)""y=C_4\cos x+C_3\sin x-\dfrac{1}{2}x^2\cos x+\dfrac{1}{2}x+\dfrac{1}{4}\cos x"

The general solution of the non homogeneous differential equation is


"y=C_5\cos x+C_3\sin x-\dfrac{1}{2}x^2\cos x+\dfrac{1}{2}x"